$9000 is invested in two accounts. Account A pays 4% annual interest and account B pay 6% annual interest. In one year a total of $383 is earned in interest. How much was invested in each account.
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y =
9000
Interest----- 383

Part I 4.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 9000 -------------1
Interest on both investments
4.00% x + 6.00% y= 383
Multiply by 100
4 x + 6 y= 38300.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -36000.00
Add this to (2)
0 x 2 y= 2300
divide by 2
y = 1150
Part I 4.00% $ 7850
Part II 6.00% $ 1150

CHECK
7850 --------- 4.00% ------- 314.00
1150 ------------- 6.00% ------- 69.00
Total -------------------- 383.00

m.ananth@hotmail.ca